if a certain type of bronze requires 0.3 kilogram of tin per 1.0 kilogram of copper how much tin must be combined with 50 kilograms of copper?
Answer by CubeyThePenguin(3113)

0.3/1 = x/50

x = (0.3)(50) = 15 kg tin